OpenStudy (anonymous):

A 24-foot beam is to be cut in two. If one portion is half the length of the other portion, what are the two lengths of the portions

OpenStudy (anonymous):

Let x equal the length of the first portion\[x+\frac{x}{2}=24, \frac{3 x}{2}=24,x=16, x/2=8 \]